References

[1] M.F. Atiyah, N.J. Hitchin and I.M. Singer, Self-duality in four-dimensional Riemannian geometry, Proc. Roy. Soc. London, A 362 (1978), 425-461.

[2] A. Boggess, CR manifolds and the tangential Cauchy-Riemann complex, CRC Press, Boca Raton, 1993.

[3] Th. Friedrich and R. Grunewald, On Einstein metrics on the twistor space of a four-dimensional Riemannian manifold, Math. Nachr. 123 (1985), 55-60.

[4] Th. Friedrich and H. Kurke, Compact four-dimensional self-dual Einstein manifolds with positive scalar curvature, Math. Nachr. 106 (1982), 271-299.

[5] G.R. Jensen and M. Rigoli, Twistor and Gauss lifts of surfaces in four-manifolds, Contemporary Math., American Math. Soc., 132 (1989), 197-232.

[6] G.R. Jensen and M. Rigoli Neutral Surfaces in Neutral Four-space, Le Mathematiche, 45 (1990), 407-443.

[7] W.F. Kamber and Ph. Tondeur, Foliations and Metrics, In: Differential Geometry, Proc. Special Year, Maryland 1981-82 (Brooks R., Gray A. and Reinhart B. L., eds.), Progress in Math. 32, Birkhäuser, Boston, 1983, 103-152.

[8] G. Legrand, Étude d’une généralisation des structures presque complexes sur les variétés différentiables, Rend. Circolo Mat. Palermo, 7 (1958), 323-354 and 8 (1959), 5-48.

[9] S. Kobayashi and K. Nomizu, Foundations of Differential Geometry I, II, Intersc. Publ., New York, 1963, 1969.

[10] P. Libermann Ch.-M. Marle, Symplectic Geometry and Analytical Mechanics, D. Reidel Publ. Comp., Dordrecht-Boston, 1987.

[11] L. Nirenberg, A complex Frobenius theorem. Seminar on Analytic Functions, Inst. for Adv. Studies, Princeton, I (1957), 172-179.

[12] N.R. O’Brian and J.H. Rawnsley, Twistor Spaces, Annals of Global Analysis and Geometry, 3 (1985), 29-58.

[13] R. Penrose, The twistor programme, Reports on Math. Physics, 12 (1977), 65-76.

[14] B.A. Rozenfeld, Multidimensional Spaces, Nauka, Moskwa, 1966. (In Russian.)

[15] I. Vaisman, On locally conformal almost Kähler manifolds, Israel J. Math., 24 (1976), 338-351.

[16] I. Vaisman, Riemannian manifolds with a parallel field of complex planes, J. of Geometry, 17 (1981), 174-192.

[17] I. Vaisman, Locally Conformal Symplectic Manifolds, International J. of Math. Sci., 8 (1985), 521-536.

[18] I. Vaisman, Symplectic Curvature Tensors, Monatshefte für Math., 100 (1985), 299-327.

[19] I. Vaisman, Symplectic Twistor Spaces, J. Geom and Physics, 3 (1986), 507-524.

[20] I. Vaisman, Symplectic Geometry and Secondary Characteristic Classes, Progress in Math., 72, Birkhäuser, Boston, 1987.

[21] I. Vaisman, On the Geometry of the Lagrangian-Grassmannian Bundles, Algebras, Groups and Geometries, 13 (1996), 323-341.

[22] I.Vaisman, On the twistor spaces of almost Hermitian manifolds, Annal. Global Anal. Geom., 16(1998), 335-356.

Department of Mathematics University of Haifa, Israel email: vaisman@math.haifa.ac.il